# Adaptive Reward Design for Reinforcement Learning in Complex Robotic Tasks

### v1

@@ Abstract
There is a surge of interest in using formal languages such as Linear Temporal Logic (LTL) and finite automata to precisely and succinctly specify complex tasks and derive reward functions for reinforcement learning (RL) in robotic applications. However, existing methods often assign sparse rewards (e.g., giving a reward of 1 only if a task is completed and 0 otherwise), necessitating extensive exploration to converge to a high-quality policy. To address this limitation, we propose a suite of reward functions that incentivize an RL agent to make measurable progress on tasks specified by LTL formulas and develop an adaptive reward shaping approach that dynamically updates these reward functions during the learning process. Experimental results on a range of RL-based robotic tasks demonstrate that the proposed approach is compatible with various RL algorithms and consistently outperforms baselines, achieving earlier convergence to better policies with higher task success rates and returns. Code is available at https://github.com/RewardShaping/AdaptiveRewardShaping.
I Introduction
@@
Complex robotic tasks require reinforcement learning (RL) frameworks capable of handling intricate task specifications. An RL agent’s behavior is guided by reward functions, which are often challenging to define manually in robotic settings due to the nuanced and hierarchical nature of robotic objectives. Alternatively, an RL agent can infer intended rewards from demonstrations [1], trajectory comparisons [2], or human instructions [3]. Recently, formal languages such as Linear Temporal Logic (LTL) and finite automata have gained traction as tools for specifying robotic tasks and deriving reward functions in RL (see the extensive list of related work in Section I-A). However, existing methods often rely on sparse rewards (e.g., assigning a reward of 1 only upon task completion and 0 otherwise), which significantly limits their efficiency for robotic applications by requiring extensive exploration to converge to effective policies. Additionally, many prior works are restricted to specific RL algorithms tailored to their reward structures, such as Q-learning for reward machines [4], modular DDPG [5], and hierarchical RL for reward machines [6], posing further challenges for deploying RL in diverse robotic systems.
@@
Reward shaping [7] is a paradigm where an agent receives some intermediate rewards as it gets closer to the goal and has shown to be helpful for RL algorithms to converge more quickly. Inspired by this idea, we develop a logic-based adaptive reward shaping approach in this work. We use the syntactically co-safe fragment of LTL to specify complex RL tasks, such as “the robotic task is to touch red and green balls in strict order without touching other colors, then touch blue balls”. We then translate a co-safe LTL task specification into a deterministic finite automaton (DFA) and design reward functions that keeps track of the task completion status (i.e., a task is completed if an accepting state of the DFA has been reached).
@@
The principle underlying our approach is that we want to assign intermediate rewards to an agent as it makes progress toward completing a task. A key challenge is how to measure the closeness to task completion. We adopt the notion of task progression for robotics defined by [8], which measures each DFA state’s distance to accepting states. The smaller the distance, the higher degree of task progression. The distance is zero when a robotic task is fully completed.
@@
Finally, we evaluate the proposed approach on various discrete and continuous robotic tasks. Computational experiments show the compatibility of our approach with a wide range of RL algorithms. Results indicate our approach generally outperforms baselines, achieving earlier convergence to better policies with higher task success rates and returns.
I-A Related Work
@@
[9] presents one of the first works applying temporal logic to reward function design, assigning reward functions based on robustness degrees of satisfying truncated LTL formulas. [10] uses a fragment of LTL for finite traces (called LTLf) to encode RL rewards. Several methods seek to learn optimal policies that maximize the probability of satisfying an LTL formula [11, 12, 5]. However, these methods assign sparse rewards for task completion and do not provide intermediate rewards for task progression.
@@
There is a line of work on reward machines (RMs), a type of finite state machine that takes labels representing environment abstractions as input and outputs reward functions. [4] shows that LTL and other regular languages can be automatically translated into RMs via the construction of DFAs. [6] describes a collection of RL methods that exploit the RM structure, including Q-learning for reward machines (QRM), counterfactual experiences for reward machines (CRM), and hierarchical RL for reward machines (HRM). These methods are augmented with potential-based reward shaping [7], where a potential function over RM states is computed to assign intermediate rewards. We adopt these methods (with reward shaping) as baselines for comparison in our experiments. As we will show in Section V, our approach generally outperforms baselines, providing more effective design of intermediate rewards for task progression.
@@
[13] proposes a new specification language that can be translated into reward functions and later applies it for compositional RL in [14]. These methods use a task monitor to track the degree of specification satisfaction and assign intermediate rewards. However, they require users to encode atomic predicates into quantitative values for reward assignment. In contrast, our approach automatically assigns intermediate rewards using DFA states’ distance to acceptance values, eliminating the need for user-provided functions.
@@
[15] presents a reward shaping framework for average-reward learning in continuing tasks. Their method automatically translates a LTL formula encoding domain knowledge into a function that provides additional reward throughout the learning process. This work has a different problem setup and thus is not directly comparable with our approach.
@@
[16] proposes an approach that decomposes an LTL mission into sub-goal-reaching tasks solved in a distributed manner. The same authors also present a model-free RL method for minimally violating an infeasible LTL specification in [17]. Both works consider the assignment of intermediate rewards, but their definition of task progression requires additional information about the environment (e.g., geometric distance from each waypoint to the destination). In contrast, we define task progression based solely on the task specification, following [8], which is a work on robotic planning with MDPs (but not RL).
II Background
II-A Reinforcement Learning
@@
II-B Co-Safe LTL Specifications
We utilize Linear Temporal Logic (LTL) [18], which is a form of modal logic that augments propositional logic with temporal operators, to specify complex tasks for the robotic agent. We focus on the syntactically co-safe LTL fragment, defined as follows.
@@ II-B Co-Safe LTL Specifications
II-C Task Progression
We adopt the notion of “task progression” introduced in [8] to measure the degree to which a robotic task defined by a co-safe LTL formula $\varphi$ is completed.
@@ II-C Task Progression
The first condition mandates $q^{\prime}\not\to^{*}q$ to ensure that there is no cycle in the DFA with a non-zero progression value, which is crucial for the convergence of infinite sums of progression [8].
@@ III Problem Formulation
IV Approach
To solve this problem, we design two reward functions that incentivize an RL agent to improve the task progression (cf. Section IV-A), and develop an adaptive reward shaping approach that dynamically updates these reward functions during the learning process (cf. Section IV-B).
IV-A Basic Reward Functions
@@
Example 4. Assuming a deterministic environment for the gridworld shown in Figure 1(a), the MDP has a discount factor of $\gamma=0.9$. We calculate the expected returns for policies from Example III using the progression reward function. ${V_{\mathsf{pg}}^{\pi_{1}}}(s_{0}^{\otimes})=0.9^{9}\approx 0.39$, ${V_{\mathsf{pg}}^{\pi_{2}}}(s_{0}^{\otimes})=0.9^{15}+0.9^{19}\approx 0.34$, and ${V_{\mathsf{pg}}^{\pi_{3}}}(s_{0}^{\otimes})=0$. Among these policies, $\pi_{1}$ yields the highest expected return, yet it fails to achieve the best possible task progression, as $b(\pi_{1})=1>b^{*}=0$. $\blacksquare$
@@
Example 5. We calculate the expected returns of policies in Example III using the hybrid reward function (with $\eta=0.1$). ${V_{\mathsf{hd}}^{\pi_{1}}}(s_{0}^{\otimes})\approx-1.15$, ${V_{\mathsf{hd}}^{\pi_{2}}}(s_{0}^{\otimes})\approx-1.33$, and ${V_{\mathsf{hd}}^{\pi_{3}}}(s_{0}^{\otimes})\approx-0.69$. Although $\pi_{3}$ yields the highest expected return, it falls short in the task progression with $b(\pi_{3})=2$. Increasing $\eta$ emphasizes penalties without altering the optimal policy in this example. Conversely, reducing $\eta$ moves closer to the progression reward function, especially when $\eta=0$. $\blacksquare$
IV-B Adaptive Reward Shaping
While reward functions defined in Section IV-A motivate an RL agent to complete a task specified by a co-safe LTL formula, Examples IV-A and IV-A show that the learned optimal policies that maximize the expected return do not achieve the best possible task progression. A potential reason is that the distance-to-acceptance function $d_{\varphi}$, as defined in Equation 1, may not precisely reflect the difficulty of activating desired DFA transitions within a specific environment. To tackle this limitation, we develop an adaptive reward shaping approach that dynamically updates distance-to-acceptance values and reward functions during the learning process.
@@ IV-B Adaptive Reward Shaping
where $b_{k}$ is the task progression of the optimal policy learned after $k\cdot N$ episodes, and $\theta$ is a hyperparameter, also used later in Equation IV-B, requiring that $\theta>1$.
Example 6. We have $d^{0}_{\varphi}(q_{0})=2$, $d^{0}_{\varphi}(q_{1})=d^{0}_{\varphi}(q_{2})=1$, $d^{0}_{\varphi}(q_{3})=15$, and $d^{0}_{\varphi}(q_{4})=0$ following Example II-C. Suppose $\pi_{1}$ is the optimal policy learned after the first $N$ episodes and thus $b_{1}=1$. Let $\theta=100$. For states in $B_{1}\cup B_{2}\cup B_{3}=\{q_{0},q_{1},q_{2},q_{3}\}$, We update their distance-to-acceptance values as follows: $d^{1}_{\varphi}(q_{1})=d^{1}_{\varphi}(q_{2})=101$, $d^{1}_{\varphi}(q_{0})=102$, and $d^{1}_{\varphi}(q_{3})=115$. For state $q_{4}\in B_{0}$, we retain its distance-to-acceptance value as $d^{1}_{\varphi}(q_{4})=0$. $\blacksquare$
@@
When $k=0$, the adaptive progression reward function ${R_{\mathsf{ap},0}^{\otimes}}$ coincides with the progression reward function ${R_{\mathsf{pg}}^{\otimes}}$ defined in Equation IV-A.
Example 7. Using the updated distance-to-acceptance values from Example IV-B, we calculate the adaptive progression rewards ${R_{\mathsf{ap},1}^{\otimes}}$ for the first round of update. For instance, we have $\rho^{1}_{\varphi}(q_{1},q_{4})=\max\{0,d^{1}_{\varphi}(q_{1})-d^{1}_{\varphi}(q_{4})\}=101$. Recall $\rho^{0}_{\varphi}(q_{1},q_{4})=1$ from Example II-C. Thus,
@@
The expected returns of policies in Example III with ${R_{\mathsf{ap},1}^{\otimes}}$ are ${V_{\mathsf{ap},1}^{\pi_{1}}}(s_{0}^{\otimes})\approx 0.39$, ${V_{\mathsf{ap},1}^{\pi_{2}}}(s_{0}^{\otimes})\approx 13.85$, and ${V_{\mathsf{ap},1}^{\pi_{3}}}(s_{0}^{\otimes})=0$. Policy $\pi_{2}$ yields the highest expected return while completing the task (i.e., $b(\pi_{2})=0$). $\blacksquare$
@@
Example 8. Let $\eta_{0}=0.1$, and $\theta_{1}=100$. The initial distance-to-acceptance values $d^{0}_{\varphi}$ are the same as in Example IV-B. Suppose the agent’s movement during the episodes follows a policy $\pi$ such that $b(\pi)=1$. Following Equation 5, we update the distance-to-acceptance values of states in $B_{1}\cup B_{2}\cup B_{3}=\{q_{0},q_{1},q_{2},q_{3}\}$ to $d^{1}_{\varphi}(q_{1})=d^{1}_{\varphi}(q_{2})=101$, $d^{1}_{\varphi}(q_{0})=102$, and $d^{1}_{\varphi}(q_{3})=115$. We compute ${R_{\mathsf{ah},1}^{\otimes}}$ with $\eta_{1}=0.001$, which yields ${V_{\mathsf{ah},1}^{\pi_{1}}}(s_{0}^{\otimes})\approx-0.52$, ${V_{\mathsf{ah},1}^{\pi_{2}}}(s_{0}^{\otimes})\approx 12.97$, and ${V_{\mathsf{ah},1}^{\pi_{3}}}(s_{0}^{\otimes})\approx-0.35$. The optimal policy $\pi_{2}$ not only yields the highest expected return but also completes the task with $b(\pi_{2})=0$. $\blacksquare$
Correctness. The correctness of the proposed adaptive reward shaping approach, as it pertains to the problem formulated in Section III, is stated below, with the proof provided in Appendix A.
@@
V Experiments
We evaluate the proposed adaptive reward shaping approach in a variety of benchmark RL domains. We describe the experimental setup including environments, RL algorithms, baselines, and evaluation metrics in Section V-A, and analyze the experimental results in Section V-B.
V-A Experimental Setup

### v2

@@ Abstract
There is a surge of interest in using formal languages such as Linear Temporal Logic (LTL) to precisely and succinctly specify complex tasks and derive reward functions for Reinforcement Learning (RL). However, existing methods often assign sparse rewards (e.g., giving a reward of 1 only if a task is completed and 0 otherwise). By providing feedback solely upon task completion, these methods fail to encourage successful subtask completion. This is particularly problematic in environments with inherent uncertainty, where task completion may be unreliable despite progress on intermediate goals. To address this limitation, we propose a suite of reward functions that incentivize an RL agent to complete a task specified by an LTL formula as much as possible, and develop an adaptive reward shaping approach that dynamically updates reward functions during the learning process. Experimental results on a range of benchmark RL environments demonstrate that the proposed approach generally outperforms baselines, achieving earlier convergence to a better policy with higher expected return and task completion rate. Code is available at https://github.com/safe-autonomy-lab/AdaptiveRewardRL.git.
1 Introduction
@@
In reinforcement learning (RL), an agent’s behavior is guided by reward functions, which are often difficult to specify manually when representing complex tasks. Alternatively, an RL agent can infer the intended reward from demonstrations Ng and Russell [2000], trajectory comparisons Wirth et al. [2017], or human instructions Fu et al. [2018]. Recent years have seen a surge of interest in using formal languages such as Linear Temporal Logic (LTL) and finite automata to specify complex tasks and derive reward functions for RL (see the extensive list of related work in Section 1.1). Nevertheless, existing methods often assign sparse rewards (e.g., giving a reward of 1 only if a task is completed and 0 otherwise). Sparse rewards may necessitate hundreds of thousands of exploratory episodes for convergence to a quality policy. Furthermore, many prior works are only compatible with specific RL algorithms tailored to their proposed reward structures, such as Q-learning for reward machines Camacho et al. [2019], modular DDPG Hasanbeig et al. [2020], and hierarchical RL for reward machines Icarte et al. [2022].
@@
Reward shaping Ng et al. [1999] is a paradigm where an agent receives some intermediate rewards as it gets closer to the goal and has shown to be helpful for RL algorithms to converge more quickly. Inspired by this idea, we develop a logic-based adaptive reward shaping approach in this work. We use the syntactically co-safe fragment of LTL to specify complex RL tasks, such as “the task is to touch red and green balls in strict order without touching other colors, then touch blue balls”. We then translate a co-safe LTL task specification into a deterministic finite automaton (DFA) and design reward functions that keeps track of the task completion status (i.e., a task is completed if an accepting state of the DFA has been reached).
@@
The principle underlying our approach is to assign intermediate rewards to an agent as it makes progress toward completing a task. A key challenge is how to measure the closeness to task completion. We adopt the notion of task progression defined by Lacerda et al. [2019], which measures each DFA state’s distance to accepting states. The smaller the distance, the higher degree of task progression. The distance is zero when the task is fully completed.
@@
Finally, we evaluate the proposed approach on various discrete and continuous RL environments. Computational experiments show the compatibility of our approach with a wide range of RL algorithms. Results indicate our approach generally outperforms baselines, achieving earlier convergence to a better policy with higher expected return and task completion rate.
1.1 Related Work
@@
Li et al. [2017] presents one of the first works applying temporal logic to reward function design, assigning reward functions based on robustness degrees of satisfying truncated LTL formulas. De Giacomo et al. [2019] uses a fragment of LTL for finite traces (called LTLf) to encode RL rewards. Several methods seek to learn optimal policies that maximize the probability of satisfying an LTL formula Hasanbeig et al. [2019], Bozkurt et al. [2020], Hasanbeig et al. [2020]. However, these methods assign sparse rewards for task completion and do not provide intermediate rewards for task progression.
@@
There is a line of work on reward machines (RMs), a type of finite state machine that takes labels representing environment abstractions as input and outputs reward functions. Camacho et al. [2019] shows that LTL and other regular languages can be automatically translated into RMs via the construction of DFAs. Icarte et al. [2022] describes a collection of RL methods that exploit the RM structure, including Q-learning for reward machines (QRM), counterfactual experiences for reward machines (CRM), and hierarchical RL for reward machines (HRM). These methods are augmented with potential-based reward shaping Ng et al. [1999], where a potential function over RM states is computed to assign intermediate rewards. We adopt these methods (with reward shaping) as baselines for comparison in our experiments. As we will show in Section 5, our approach generally outperforms baselines, providing more effective design of intermediate rewards for task progression.
@@
Jothimurugan et al. [2019] proposes a new specification language that can be translated into reward functions and later applies it for compositional RL in Jothimurugan et al. [2021]. These methods use a task monitor to track the degree of specification satisfaction and assign intermediate rewards. However, they require users to encode atomic predicates into quantitative values for reward assignment. In contrast, our approach automatically assigns intermediate rewards using DFA states’ distance to acceptance values, eliminating the need for user-provided functions.
@@
Jiang et al. [2021] presents a reward shaping framework for average-reward learning in continuing tasks. Their method automatically translates a LTL formula encoding domain knowledge into a function that provides additional reward throughout the learning process. This work has a different problem setup and thus is not directly comparable with our approach.
@@
Cai et al. [2023a] proposes an approach that decomposes an LTL mission into sub-goal-reaching tasks solved in a distributed manner. The same authors also present a model-free RL method for minimally violating an infeasible LTL specification in Cai et al. [2023b]. Both works consider the assignment of intermediate rewards, but their definition of task progression requires additional information about the environment (e.g., geometric distance from each waypoint to the destination). In contrast, we define task progression based solely on the task specification, following Lacerda et al. [2019], which is a work on robotic planning with MDPs (but not RL).
2 Background
2.1 Reinforcement Learning
@@
2.2 Co-Safe LTL Specifications
We utilize Linear Temporal Logic (LTL) Pnueli [1981], which is a form of modal logic that augments propositional logic with temporal operators, to specify complex tasks for the robotic agent. We focus on the syntactically co-safe LTL fragment, defined as follows.
@@ 2.2 Co-Safe LTL Specifications
2.3 Task Progression
We adopt the notion of “task progression” introduced in Lacerda et al. [2019] to measure the degree to which a robotic task defined by a co-safe LTL formula $\varphi$ is completed.
@@ 2.3 Task Progression
The first condition mandates $q^{\prime}\not\to^{*}q$ to ensure that there is no cycle in the DFA with a non-zero progression value, which is crucial for the convergence of infinite sums of progression Lacerda et al. [2019].
@@ 3 Problem Formulation
4 Approach
To solve this problem, we design two reward functions that incentivize an RL agent to improve the task progression (cf. Section 4.1), and develop an adaptive reward shaping approach that dynamically updates these reward functions during the learning process (cf. Section 4.2).
4.1 Basic Reward Functions
@@
Example 4. Assuming a deterministic environment for the gridworld shown in Figure 1(a), the MDP has a discount factor of $\gamma=0.9$. We calculate the expected returns for policies from Example 3 using the progression reward function. ${V_{\mathsf{pg}}^{\pi_{1}}}(s_{0}^{\otimes})=0.9^{9}\approx 0.39$, ${V_{\mathsf{pg}}^{\pi_{2}}}(s_{0}^{\otimes})=0.9^{15}+0.9^{19}\approx 0.34$, and ${V_{\mathsf{pg}}^{\pi_{3}}}(s_{0}^{\otimes})=0$. Among these policies, $\pi_{1}$ yields the highest expected return, yet it fails to achieve the best possible task progression, as $b(\pi_{1})=1>b^{*}=0$. $\blacksquare$
@@
Example 5. We calculate the expected returns of policies in Example 3 using the hybrid reward function (with $\eta=0.1$). ${V_{\mathsf{hd}}^{\pi_{1}}}(s_{0}^{\otimes})\approx-1.15$, ${V_{\mathsf{hd}}^{\pi_{2}}}(s_{0}^{\otimes})\approx-1.33$, and ${V_{\mathsf{hd}}^{\pi_{3}}}(s_{0}^{\otimes})\approx-0.69$. Although $\pi_{3}$ yields the highest expected return, it falls short in the task progression with $b(\pi_{3})=2$. Increasing $\eta$ emphasizes penalties without altering the optimal policy in this example. Conversely, reducing $\eta$ moves closer to the progression reward function, especially when $\eta=0$. $\blacksquare$
4.2 Adaptive Reward Shaping
While reward functions defined in Section 4.1 motivate an RL agent to complete a task specified by a co-safe LTL formula, Examples 4.1 and 4.1 show that the learned optimal policies that maximize the expected return do not achieve the best possible task progression. A potential reason is that the distance-to-acceptance function $d_{\varphi}$, as defined in Equation 1, may not precisely reflect the difficulty of activating desired DFA transitions within a specific environment. To tackle this limitation, we develop an adaptive reward shaping approach that dynamically updates distance-to-acceptance values and reward functions during the learning process.
@@ 4.2 Adaptive Reward Shaping
where $b_{k}$ is the task progression of the optimal policy learned after $k\cdot N$ episodes, and $\theta$ is a hyperparameter, also used later in Equation 4.2, requiring that $\theta>1$.
Example 6. We have $d^{0}_{\varphi}(q_{0})=2$, $d^{0}_{\varphi}(q_{1})=d^{0}_{\varphi}(q_{2})=1$, $d^{0}_{\varphi}(q_{3})=15$, and $d^{0}_{\varphi}(q_{4})=0$ following Example 2.3. Suppose $\pi_{1}$ is the optimal policy learned after the first $N$ episodes and thus $b_{1}=1$. Let $\theta=100$. For states in $B_{1}\cup B_{2}\cup B_{3}=\{q_{0},q_{1},q_{2},q_{3}\}$, We update their distance-to-acceptance values as follows: $d^{1}_{\varphi}(q_{1})=d^{1}_{\varphi}(q_{2})=101$, $d^{1}_{\varphi}(q_{0})=102$, and $d^{1}_{\varphi}(q_{3})=115$. For state $q_{4}\in B_{0}$, we retain its distance-to-acceptance value as $d^{1}_{\varphi}(q_{4})=0$. $\blacksquare$
@@
When $k=0$, the adaptive progression reward function ${R_{\mathsf{ap},0}^{\otimes}}$ coincides with the progression reward function ${R_{\mathsf{pg}}^{\otimes}}$ defined in Equation 4.1.
Example 7. Using the updated distance-to-acceptance values from Example 4.2, we calculate the adaptive progression rewards ${R_{\mathsf{ap},1}^{\otimes}}$ for the first round of update. For instance, we have $\rho^{1}_{\varphi}(q_{1},q_{4})=\max\{0,d^{1}_{\varphi}(q_{1})-d^{1}_{\varphi}(q_{4})\}=101$. Recall $\rho^{0}_{\varphi}(q_{1},q_{4})=1$ from Example 2.3. Thus,
@@
The expected returns of policies in Example 3 with ${R_{\mathsf{ap},1}^{\otimes}}$ are ${V_{\mathsf{ap},1}^{\pi_{1}}}(s_{0}^{\otimes})\approx 0.39$, ${V_{\mathsf{ap},1}^{\pi_{2}}}(s_{0}^{\otimes})\approx 13.85$, and ${V_{\mathsf{ap},1}^{\pi_{3}}}(s_{0}^{\otimes})=0$. Policy $\pi_{2}$ yields the highest expected return while completing the task (i.e., $b(\pi_{2})=0$). $\blacksquare$
@@
Example 8. Let $\eta_{0}=0.1$, and $\theta_{1}=100$. The initial distance-to-acceptance values $d^{0}_{\varphi}$ are the same as in Example 4.2. Suppose the agent’s movement during the episodes follows a policy $\pi$ such that $b(\pi)=1$. Following Equation 5, we update the distance-to-acceptance values of states in $B_{1}\cup B_{2}\cup B_{3}=\{q_{0},q_{1},q_{2},q_{3}\}$ to $d^{1}_{\varphi}(q_{1})=d^{1}_{\varphi}(q_{2})=101$, $d^{1}_{\varphi}(q_{0})=102$, and $d^{1}_{\varphi}(q_{3})=115$. We compute ${R_{\mathsf{ah},1}^{\otimes}}$ with $\eta_{1}=0.001$, which yields ${V_{\mathsf{ah},1}^{\pi_{1}}}(s_{0}^{\otimes})\approx-0.52$, ${V_{\mathsf{ah},1}^{\pi_{2}}}(s_{0}^{\otimes})\approx 12.97$, and ${V_{\mathsf{ah},1}^{\pi_{3}}}(s_{0}^{\otimes})\approx-0.35$. The optimal policy $\pi_{2}$ not only yields the highest expected return but also completes the task with $b(\pi_{2})=0$. $\blacksquare$
Correctness. The correctness of the proposed adaptive reward shaping approach, as it pertains to the problem formulated in Section 3, is stated below, with the proof provided in Appendix A.
@@
5 Experiments
We evaluate the proposed adaptive reward shaping approach in a variety of benchmark RL domains. We describe the experimental setup including environments, RL algorithms, baselines, and evaluation metrics in Section 5.1, and analyze the experimental results in Section 5.2.
5.1 Experimental Setup